OpenStudy (lovelyanna):

Which intervals are most appropriate for constructing a frequency table for the data set? 65 73 80 56 42 47 67 88 53 71 75 49 48 65 63 A. 40–55, 55–70, 70–85, 85–100 B. 40–54, 55–69, 70–84, 85–99 C. 40–59, 60–74, 75–85, 86–95 D. 40–59, 60–69, 70–79, 80–89

OpenStudy (lovelyanna):

Is it D?

OpenStudy (math&ing001):

Let me refresh my memory on this :)

OpenStudy (lovelyanna):

Ok :)

OpenStudy (math&ing001):

Ok, so apparently this is how you do it : 1- Highest value = 88 Lowest value = 42 Difference = 88 - 42 = 46 Let the width be 15. 46/15=3 The classes would be 40-55 55-70 70-85 85-100 It can't be D because the intervals are not equal : 59-40=19 and 69-60=9

OpenStudy (lovelyanna):

Mhm.. I see how it's done, my mistake, sorry :P So the answer is A

OpenStudy (math&ing001):

Yep ! See : http://www.mathsteacher.com.au/year8/ch17_stat/03_freq/freq.htm if you want more info.

OpenStudy (lovelyanna):

Thanks! :)

OpenStudy (math&ing001):

You're always welcome :3